# Lindblad-like quantum tomography for non-Markovian quantum dynamical maps

### v1

@@ I Introduction
However, noise in real QIPs does not always fall in this category, and temporal correlations and even non-Markovianity can play an important role, as alluded above in connection to QES, QEM and QEC. Hence, it would be desirable to extend the Lindblad learning to encompass non-Markovian noise scenarios, such that the quantum dynamical maps are no longer a semigroup, nor can they be divided into the composition of sequential CP channels at any intermediate time [75, 76]. In general, any quantum evolution that results from the coupling of a quantum system to a larger environment, or to a set of noisy controls modeled by stochastic processes, can be expressed in terms of a time-local master equation by using a time-convolutionless formulation [91] of the Nakajima-Zwanzig integro-differential equation [92, 93, 74]. These time-local master equations generalize the aforementioned Lindblad master equation [78, 79], and can be expressed in a canonical form that connects directly with the degree of non-Markovianity [94]. In essence, the characterization of these time-local master equations would require a time-dependent parametrization of the Hamiltonian, jump operators and dissipation rates, which can then be incorporated into a maximum-likelihood estimation that parallels the Markovian Lindblad limit [87, 88, 89, 90]. In this work, we call this QCVV technique Lindblad-like quantum tomography ($\ell$QT), and develop it in the simplest possible scenario: the dephasing dynamics of a single qubit. We present a detailed comparative study of this QCVV technique, considering both a frequentist and a Bayesian approach for the statistical inference. We consider minimal dephasing models, both semi-classical and fully quantum-mechanical, in which the temporal correlations and degree of non-Markovianity can be independently controlled. By making a careful connection to the the theory of asymptotic inference and Bayesian estimation, we quantify both the accuracy and precision of $\ell$QT. We discuss how the amount of temporal correlations and the degree of non-Markovianity can play a key role in deciding which of the two approaches is preferable when learning the non-Lindblad qubit dephasing.
@@
This article is organized as follows. In Sec. II we review the techniques of Lindblad quantum tomography. Sec. III presents $\ell$QT, a generalization of Lindblad learning that allows us to characterize non-Markovian dephasing noise. In Secs. III.1 and III.2 two approaches to $\ell$QT are presented: a frequentist and a Bayesian approach. Both are compared in a performance analysis in Sec. III.3, where we also study how measurement times should be selected to reduce the number of necessary measurements and the error in the estimation of noise parameters. We conclude in Sec. IV.
@@
II Learning time-local master equations by maximum-likelihood estimation
@@
LQT makes use of a total of $N_{\rm shot}$ measurement shots, also known as trials in the context of statistics, which will be distributed among the different instants of time $N_{\rm shot}=\sum_{i}N_{i}$. These will themselves be also distributed among the different initial states and measurement basis $N_{i}=\sum_{s,b}N_{s,i,b}$. In the experiment, one would count the number of times $N_{s,i,b,m_{b}}$ that the $m_{b}$ outcome is obtained for each of the configurations, such that $N_{s,i,b}=\sum_{m_{b}}N_{s,i,b,m_{b}}$. This provides a data set $\mathbb{D}=\{N_{s,i,b,m_{b}}\}$ that can be understood as a random sample of the corresponding random variable $\tilde{N}_{s,i,b,m_{b}}$ obtained from $N_{\rm shot}$ experimental measurements. Using the convention of App. A, we use tildes to refer to stochastic variables and stochastic processes, which would be described by an underlying probability distribution. In this case, the joint probability distribution can be defined from the postulates of quantum mechanics, taking into account Born’s rule
$p_{s,i,\mu}=\mathrm{Tr}\left\{M_{\mu}\mathcal{E}_{t_{i},t_{0}}(\rho_{0,s})\right\}=:p_{s,i,b}(m_{b}),$ (3)
where $\mathcal{E}_{t,t_{0}}\in\mathsf{C}(\mathcal{H}_{\rm S})$ is a one-parameter family of completely-positive trace-preserving (CPTP) channels [64] describing the actual time evolution of the noisy quantum system. In the right-hand side of Eq. (3), we use a more familiar notation in statistics [96], corresponding to the discrete probability density function (pdf) of a Bernoulli random variable with two outcomes $p_{s,i,b}(0),p_{s,i,b}(1)=1-p_{s,i,b}(0)$, such that the lower multi-index specifies the whole history of the random process, including the initial preparation step $s\in\mathbb{S}_{0}$, the evolution time $i\in\mathbb{I}_{t}$, and the final measurement basis $b\in\mathbb{M}_{b}$. The data set $\mathbb{D}$ is thus a random sample of the joint multinomial pdf for the sum of Bernoulli random variables, each of which corresponds to a separate experimental run, namely
$p_{\tilde{N}}(\mathbb{D})=\frac{N_{\rm s}!}{\prod_{s,i,\mu}{N_{s,i,\mu}!}}\prod_{s,i,b,m_{b}}\big{(}p_{s,i,b}(m_{b})\big{)}^{N_{s,i,b,m_{b}}}.$ (4)
The data set $\mathbb{D}$ can be used to estimate the Hamiltonian $H$ and Lindblad matrix $G$ by maximizing the likelihood function, which is defined by the above joint pdf with individual probabilities (3) approximated by the Lindbladian (1), namely
$\mathsf{L}(H,G|\mathbb{D})=p_{\tilde{N}}(\mathbb{D})\Bigl{|}_{p^{\phantom{L}}_{s,i,b}(m_{b})\mapsto p^{\rm L}_{s,i,b}(m_{b})=\mathrm{Tr}\{M_{\mu}\mathord{\rm e}^{(t_{i}-t_{0})\mathcal{L}_{H,G}}(\rho_{0,s})\}}.$ (5)
The larger this likelihood is for a given pair $H,G$, the better the Lindbladian description approximates the observed data [96]. LQT works by rescaling the negative logarithm of this likelihood, converting the optimization problem into a non-linear minimization of the Lindbladian cost function
$\mathsf{C}_{\rm L}(H,G)=-\sum_{s,i,b.m_{b}}N_{s,i,b,m_{b}}\log p^{\rm L}_{s,i,b}(m_{b}).$ (6)
@@
By minimizing this non-convex estimator or, instead, a convex approximation based on linearization and/or compressed sensing [90], LQT provides an estimate of the generators $\hat{H},\hat{G}$ that yield the best match with the observed data, where we will use hats to refer to estimated quantities. We note that this minimization is subject to constraints on Hamiltonian Hermicity and Lindblad matrix semidefinite positiveness.
According to a definition of non-Markovianity for quantum evolution [97, 51], the Lindbladian quantum dynamical map that results from the estimation is time-homogeneous $\mathcal{E}^{\rm L}_{t,t_{0}}(\rho_{0})=\mathcal{E}^{\rm L}_{t-t_{0}}(\rho_{0})={\rm exp}\{(t-t_{0})\mathcal{L}(\hat{H},\hat{G})\}(\rho_{0})$ and thus fulfills an additional CP-divisibility condition
$\mathcal{E}^{\rm L}_{t,t_{0}}=\mathcal{E}^{\rm L}_{t,t^{\prime}}\circ\mathcal{E}^{\rm L}_{t^{\prime},t_{0}}:\,\,\,\mathcal{E}^{\rm L}_{t,t^{\prime}}\in\mathsf{C}(\mathcal{H}_{\rm S}),\,\forall t^{\prime}\in[t_{0},t].$ (7)
Being a semigroup, the Lindbladian quantum dynamical map is Markovian [80, 78, 79], and always fulfills the above CP-divisibility. However, in many situations of relevance for current quantum information processors (QIPs), this is not the case, and non-Markovian effects arise which, loosely speaking, result from the noise memory on the system dynamics.
@@
In this work, we describe our first steps in the development of a learning procedure for non-Markovian quantum dynamical maps that supersedes the above LQT. In particular, we consider the statistical inference of the generators of dephasing quantum dynamical maps that need not fulfill Eq. (7). These dephasing maps $\rho(t)=\mathcal{E}^{\rm TL}_{t,t_{0}}(\rho_{0})$ no longer have a Lindbladian generator (1), but are instead governed by a time-local master equation that, when expressed in a canonical form [94], reads $\dot{\rho}=-\mathord{\rm i}[H(t),\rho]+\mathcal{D}_{\rm TCL}(\rho)$ with
$\mathcal{D}_{\rm TCL}(\rho)=\sum_{n}{\gamma}_{n}(t)\big{(}L_{n}(t)\rho L^{\dagger}_{n}(t)-\textstyle\frac{1}{2}\{L^{\dagger}_{n}(t)L_{n}(t),\rho\}\!\big{)}\!.$ (8)
@@
Here, the Hamiltonian $H(t)$, as well as the dissipative rates and jump operators ${\gamma}_{n}(t),L_{n}(t)$, can be time dependent. It is important to note that the ’rates’ are no longer required to be positive semidefinite. The possibility of encountering negative rates is directly linked with the non-Markovianity of the quantum evolution [51, 52, 75]. This time-local master equation can always be written in the form of Eq. (1) by letting $H\mapsto H(t)$ and $G\mapsto G(t)$, such that the corresponding quantum dynamical map will depend on the history of the time-dependent generators $\mathcal{E}^{\rm TL}_{t,t_{0}}=\mathcal{E}^{\rm TL}_{t,t_{0}}(\{H(t^{\prime}),G(t^{\prime})\}_{t^{\prime}})$ for all $t^{\prime}\in[t,t_{0}]$. We thus formulate a Lindblad-like quantum tomography ($\ell$QT) by upgrading the LQT cost function (6) to a time-local one that can encompass non-Markovian effects
@@
Therefore, we see that in addition to the time dependences, the dissipation matrix is no longer required to be semi-positive definite, but only Hermitian, and can thus support negative decay rates and incorporate non-Markovian effects.

### v2

@@ Introduction
However, noise in real QIPs does not always fall in this category, and temporal correlations and even non-Markovianity can play an important role, as alluded above in connection to QES, QEM and QEC. Hence, it would be desirable to extend the Lindblad learning to encompass non-Markovian noise scenarios, such that the quantum dynamical maps are no longer a semigroup, nor can they be divided into the composition of sequential CP channels at any intermediate time [75, 76]. Several characterization techniques for related problems have been developed in recent years. Quantum noise spectroscopy protocols [91, 92] have evolved into characterization tools of non-Markovian systems, as in Ref. [93], where the authors develop a filter-function formalism based on frames. A general non-Markovian quantum process framework was proposed by Ref. [94] allowing to characterize non-Markovian quantum processes using process tensor tomography (PTT). This later led to proposals of quantum non-Markovian process tomography [95]. Reference [96] introduced into this PTT frameworks the possibility of memory effects within the system control itself and from undesirable interactions between the control and the environment. This contrasts Ref. [94] where perfect gates were assumed. Reference [97] characterizes certain type of non-Markovian errors in quantum gates and [98] proposes also characterization of quantum gates but with a compressed approach that reduces the resources needed in comparison to, for instance, Ref. [96].
@@
In general, any quantum evolution that results from the coupling of a quantum system to a larger environment, or to a set of noisy controls modeled by stochastic processes, can be expressed in terms of a time-local master equation by using a time-convolutionless formulation [99] of the Nakajima-Zwanzig integro-differential equation [100, 101, 74]. These time-local master equations generalize the aforementioned Lindblad master equation [78, 79], and can be expressed in a canonical form that connects directly with the degree of non-Markovianity [102]. In essence, the characterization of these time-local master equations would require a time-dependent parametrization of the Hamiltonian, jump operators and dissipation rates, which can then be incorporated into a maximum-likelihood estimation that parallels the Markovian Lindblad limit [87, 88, 89, 90]. In this work, we call this QCVV technique Lindblad-like quantum tomography ($\ell$QT), and develop it in the simplest possible scenario: the dephasing dynamics of a single qubit. We present a detailed comparative study of this QCVV technique, considering both a frequentist and a Bayesian approach for the statistical inference. We consider minimal dephasing models, both semi-classical and fully quantum-mechanical, in which the temporal correlations and degree of non-Markovianity can be independently controlled. By making a careful connection to the the theory of asymptotic inference and Bayesian estimation, we quantify both the accuracy and precision of $\ell$QT. We discuss how the amount of temporal correlations and the degree of non-Markovianity can play a key role in deciding which of the two approaches is preferable when learning the non-Lindblad qubit dephasing.
@@
This article is organized as follows. In subsection “Learning time-local master equations by maximum-likelihood estimation” we review the techniques of Lindblad quantum tomography and present $\ell$QT, a generalization of Lindblad learning that allows us to characterize non-Markovian noise. In subsections “Frequentist approach to non-Markovian inference” and “Bayesian approach to non-Markovian inference” two approaches to $\ell$QT are presented. Subsection “Lindblad-like quantum tomography for non-Markovian dephasing” presents $\ell$QT applied to dephasing noise. The frequentist and Bayesian approaches are later compared in a performance analysis in subsections “Markovian semi-classical dephasing” and “Non-Markovian quantum dephasing”, where we also study how measurement times should be selected to reduce the number of necessary measurements and the error in the estimation of noise parameters.
Results
@@
Learning time-local master equations by maximum-likelihood estimation
@@
In the general case, LQT involves preparing an informationally complete set of initial states $s\in\mathbb{S}_{0}$, allowing the system to evolve over a set of times $i\in\mathbb{I}_{t}$, and performing measurements in different basis $b\in\mathbb{M}_{b}$, with corresponding outcomes $m_{b}\in\mathbb{M}_{m_{b}}$ (or more generally using a POVM). These independent configurations, consisting of initial states, evolution times, and measurement outcomes, provide the necessary data to estimate the Hamiltonian, dissipation rates, and jump operators, ensuring a complete reconstruction of the Lindbladian dynamics. LQT makes use of a total of $N_{\rm shot}$ measurement shots, also known as trials in the context of statistics, which will be distributed among the different initial states, instants of time and measurement basis $N_{\rm shot}=\sum_{s,i,b}N_{s,i,b}$. Therefore, the total number of measurement shots is a function of number of initial states, measurement basis, measurement outcomes and measurement times, expressed as $N_{\rm shot}(|\mathbb{S}_{0}|,|\mathbb{M}_{b}|,|\mathbb{M}_{m_{b}}|,|\mathbb{I}_{t}|)$, where $|\mathbb{A}|$ denotes the cardinality of the set $\mathbb{A}$, i.e., the number of elements in the set. In the experiment, one would count the number of times $N_{s,i,b,m_{b}}$ that the $m_{b}$ outcome is obtained for each of the configurations, such that $N_{s,i,b}=\sum_{m_{b}}N_{s,i,b,m_{b}}$. This provides a data set $\mathbb{D}=\{N_{s,i,b,m_{b}}\}$ that can be understood as a random sample of the corresponding random variable $\tilde{N}_{s,i,b,m_{b}}$ obtained from $N_{\rm shot}$ experimental measurements. We use tildes to refer to stochastic variables. In this case, the probability of obtaining the outcome $m_{b}$ when measuring in basis $b$ is given by $p_{s,i,b}(m_{b})=\mathrm{Tr}\left\{M_{\mu}\mathcal{E}_{t_{i},t_{0}}(\rho_{0,s})\right\}$, where $M_{\mu}$ represents the corresponding projector of the combination $(b,m_{b})$. $\mathcal{E}_{t,t_{0}}\in\mathsf{C}(\mathcal{H}_{\rm S})$ is a one-parameter family of completely-positive trace-preserving (CPTP) channels [64] describing the actual time evolution of the noisy quantum system. Each measurement configuration gives rise to an independent binomial distribution with $N_{s,i,b}$ trials.
The data set $\mathbb{D}$ can be used to estimate the Hamiltonian $H$ and Lindblad matrix $G$ by maximizing the likelihood function, which is defined as the probability distribution of the combination of all the independent binomials $p_{s,i,b}(m_{b})$. These binomials can be approximated by the Lindbladian of Eq. (1), namely $p^{\phantom{L}}_{s,i,b}(m_{b})\mapsto p^{\rm L}_{s,i,b}(m_{b})=\mathrm{Tr}\{M_{\mu}\mathord{\rm e}^{(t_{i}-t_{0})\mathcal{L}_{H,G}}(\rho_{0,s})\}$. The larger this likelihood function is for a given pair $H,G$, the better the Lindbladian description approximates the observed data [104]. Taking the negative logarithm of the likelihood function we obtain a cost function
$\mathsf{C}_{\rm L}(H,G)=-\sum_{s,i,b.m_{b}}N_{s,i,b,m_{b}}\log p^{\rm L}_{s,i,b}(m_{b}),$ (3)
@@
with the minimum located in the same place as the maximum of the original likelihood. The optimization problem is thus converted into a non-linear minimization of this Lindbladian cost function. By minimizing this non-convex estimator or, instead, a convex approximation based on linearization and/or compressed sensing [90], LQT provides an estimate of the generators $\hat{H},\hat{G}$ that yield the best match with the observed data, where we will use hats to refer to estimated quantities. We note that this minimization is subject to constraints on Hamiltonian hermicity and Lindblad matrix semidefinite positiveness.
@@
In this work, we describe our first steps in the development of a learning procedure for non-Markovian quantum dynamical maps that supersedes the above LQT. In particular, we consider the statistical inference of the generators of quantum dynamical maps that need not fulfill CP-divisibility, the property of a quantum dynamical map where the evolution between any two intermediate times can be described by a completely positive map [105, 51]. These maps $\rho(t)=\mathcal{E}^{\rm TL}_{t,t_{0}}(\rho_{0})$ no longer have the Lindbladian generator of Eq. (1), but are instead governed by a time-local master equation that, when expressed in a canonical form [102], reads $\dot{\rho}=-\mathord{\rm i}[H(t),\rho]+\mathcal{D}_{\rm TL}(\rho)$ with
$\mathcal{D}_{\rm TL}(\rho)=\sum_{n}{\gamma}_{n}(t)\big{(}L_{n}(t)\rho L^{\dagger}_{n}(t)-\textstyle\frac{1}{2}\{L^{\dagger}_{n}(t)L_{n}(t),\rho\}\!\big{)}\!.$ (4)
@@
Here, the Hamiltonian $H(t)$, as well as the dissipative rates and jump operators ${\gamma}_{n}(t),L_{n}(t)$, can be time dependent. It is important to note that the ‘rates’ are no longer required to be positive semidefinite. The possibility of encountering negative rates is directly linked with the non-Markovianity of the quantum evolution [51, 52, 75]. This time-local master equation can always be written in the form of Eq. (1) by letting $H\mapsto H(t)$ and $G\mapsto G(t)$, such that the corresponding quantum dynamical map will depend on the history of the time-dependent generators $\mathcal{E}^{\rm TL}_{t,t_{0}}=\mathcal{E}^{\rm TL}_{t,t_{0}}(\{H(t^{\prime}),G(t^{\prime})\})$ for all $t^{\prime}\in[t,t_{0}]$. Here, the $t^{\prime}$ argument reflects the fact that now the dynamical map between initial time $t_{0}$ and time $t$ depends on all intermediate times. We thus formulate a Lindblad-like quantum tomography ($\ell$QT) by upgrading the LQT cost function in Eq. (3) to a time-local one that can encompass non-Markovian effects
@@
subject to ${H}(t^{\prime})\in\mathsf{Herm}(\mathcal{H}_{\rm S}),G(t^{\prime})\in\mathsf{Herm}(\mathbb{C}^{d^{2}-1})$. Therefore, we see that in addition to the time dependences, the dissipation matrix is no longer required to be semi-positive definite, but only Hermitian, and can thus support negative decay rates and incorporate non-Markovian effects.